OpenStudy (anonymous):

Use the quadratic formula (ax^2+bx+c) to solve the following : 3x^2-x+2=0

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

maybe it's saying that you should use the formula: for ax^2+bx+c , you will get x values by \[x \ =\ \frac{-b \pm sqrt{(b^2-4ac)}} {2a}\]

OpenStudy (anonymous):

im getting : x1= 1 + square root of -23 / 6 = 1/6 + 1/6 square root of 23

OpenStudy (anonymous):